OpenStudy (anonymous):

(3x+3)/(x-1)=(3x+2)/(x+4)

OpenStudy (anonymous):

get common denominators

OpenStudy (anonymous):

ok. but there is where I keep getting it wrong. Do you multiply left with x+4 and the right side with x+1?

OpenStudy (anonymous):

yupp!

OpenStudy (anonymous):

3x^2+12x+3x=3x^2-3x+2x-2 3x^2+15x+12=3x^2-x-2 3x^2+15x+12-3x^2+x+2 16x/16+14/16=7/8